# A Spectral Element Enrichment Wall-Model

### v2

@@ 1 Introduction
This study seeks to make use of the high-order polynomial solution of SEM to develop a different kind of wall-model better suited for high-order methods than traditional models. SEM is our numerical method of choice for LES because of the low dissipation and dispersion errors for its cost [9] and its computational efficiency and scalability [10]. SEM has been shown to be more efficient than low-order methods at resolving the small scales of turbulence [9] [11] [12]. The wall-model builds upon prior work on the development of so-called enriched basis turbulence wall models in the context of DG [8] [13] and the continuous finite element (FEM) [14] [15] [16]. These methods leverage the polynomial solution and augment it by considering the empirical wall function as part of the solution representation. Doing so prevents spurious oscillations in the near-wall region due to insufficient polynomial resolution of the boundary layer and improves the solution representation by allowing the degrees of freedom to represent the flow fluctuations instead of the mean flow behavior. This work serves to develop the framework for enrichment wall-models in SEM so that SEM can be applied to a wider range of problems and the enrichment framework can be further leveraged for other applications.
@@ 2 Governing Equations
where $\bm{u}$ is the velocity vector, $t$ corresponds to time, $p$ is the pressure, and $\nu$ is kinematic viscosity of the fluid. The equations are solved over a computational domain $\Omega$ with boundary $\partial\Omega$. In this work, we consider implicit LES (iLES), where the numerical dissipation of the scheme is used to represent the small scales, instead of an explicit subgrid-scale turbulence model. iLES has seen success, particularly in the context of high-order methods [18] [19] [20] [21].
@@ 3 Spectral Element Method
Assuming $\bm{u}^{n-1}$ is $C^{0}$ continuous, $\bm{u}^{n}$, will be $C^{0}$ continuous if $\Delta\bm{u}^{n-1}$ is also $C^{0}$ continuous. Hence, because the operators in Eq. (21)-(22) enforce $C^{0}$ continuity, $\bm{u}^{n}$ will maintain the appropriate continuity at all timesteps.
The computational cost of SEM is a function of the number of elements, $N$ and the polynomial order $P$. Each element contains $(P+1)^{d}$ degrees of freedom per element, where $d$ is the dimension of the problem. The elements have $2d$ faces and $(P+1)^{d-1}$ degrees of freedom per face. Accounting for the co-located nodes that are equated, the global system has $NP^{d}$ degrees of freedom. Since the choice of quadrature rule results in discrete orthogonality, evaluating each entry of $\mathcal{M}$, $\mathcal{D}$, and $\mathcal{K}$ requires $O(1)$ operations. To avoid aliasing errors, the entries of $\mathcal{C}$ are overintegrated with a cost of $O((1.5(P+1))^{d})$. The global system is solved with efficient multigrid solvers [29].
@@ 4 Enrichment Method
The enriched SEM begins by putting the governing equations into weighted residual form as shown in Eqs. 3-4. However, for the enriched SEM, the solution is $(\bm{u},p)\in X_{b,\psi}^{P}(\Omega)\times Y^{P}(\Omega)$, where $X_{\psi}\subset H^{1}(\Omega)$ such that $X_{\psi}^{P}=\{u\,|\,u=u_{p}+\psi,u_{p}\in X^{P}(\Omega_{\psi})\cap X^{P}(\Omega_{p}),\psi\in H^{1}(\Omega_{p})\cap H^{1}(\Omega_{\psi}),\psi(\bm{x})=0\,\forall\bm{x}\in\Omega_{p}\}$, where $\Omega_{\psi}$ and $\Omega_{p}$ are the enriched and unenriched parts of the domain respectively, such that $\Omega_{\psi}\cup\Omega_{p}=\Omega$ and $\Omega_{\psi}$ and $\Omega_{p}$ do not overlap. $X_{b,\psi}$ is the subset of $X_{\psi}$ such that Dirichlet boundary conditions are satisfied on the boundary $\partial\Omega$. As is done in the traditional SEM, the weighted residual form is integrated by parts to give Eqs. 5-6. From here, the enriched solution representation from Eq. 24 is substituted into the weighted residual form:
@@
For the application to WMLES, we assume the enrichment function is constant within a timestep and is adapted explicitly using matching points, which will be detailed in Sec. 5. Hence, the time derivative of the enrichment function is put to the right hand side of Eq. (28). With the governing equations in this form, the left-hand side and the first three terms on the right-hand side are the same as in the traditional SEM and the last five terms on the right-hand side are additions from the enrichment [31]. This separability shows that the presented enrichment method can be implemented with the addition of associated terms into an existing solver for ease of implementation.
@@
Because the enrichment function is constant within each timestep, the change in the enriched solution is equal to the change in the polynomial component: $\Delta\underline{\bm{u}}^{e,n}=\Delta\underline{\bm{u}}^{e,n}_{p}$. Hence, in both enriched and unenriched elements, the change in the solution is only due to changes in the polynomial part. Further, let node $I$ be a node at $\bm{x}_{I}$ on the edge between elements $e_{\psi}$ and $e_{p}$ which are in the enriched and unenriched sections of the domain, respectively. Assuming that at timestep $n-1$, the solution is $C^{0}$ continuous then
$\underline{{\bm{u}}}^{e_{\psi},n-1}_{I}=\underline{{\bm{u}}}^{e_{\psi},n-1}_{I,p}+\bm{\psi}^{e_{\psi}}(\bm{x}_{I})=\underline{{\bm{u}}}^{e_{p},n-1}_{I}.$ (42)
As previously discussed, at timestep $n$, the change in the solution is only in the polynomial parts, hence at the boundary between the enriched and unenriched sections of the domain, the change in polynomial coefficients are equal:
$\underline{{\bm{u}}}^{e_{\psi},n}_{I}$ $=\underline{{\bm{u}}}^{e_{\psi},n-1}_{I,p}+\Delta\underline{{\bm{u}}}^{e_{\psi},n}_{I,p}+\bm{\psi}^{e_{\psi}}(\bm{x}_{I})=\underline{{\bm{u}}}^{e_{p},n-1}_{I}+\Delta\underline{{\bm{u}}}^{e_{p},n}_{I},$ (43)
$\Delta\underline{{\bm{u}}}^{e_{\psi},n}_{I,p}$ $=\Delta\underline{{\bm{u}}}^{e_{p},n}_{I}.$ (44)
Using the fact that the change in polynomial coefficients are equal at all element boundaries, the global system of equations can be constructed by equating $\Delta\underline{\bm{u}}_{p}$ between elements using the standard SEM scatter-gather operations [27]:
@@
The computational cost of the enrichment method is the same as standard SEM for the polynomial terms. As previously discussed, the enrichment terms require more accurate quadrature rules to integrate. Let $Q_{\psi}$ be the number of quadrature points needed to integrate the enrichment function. Then, the entries of $\mathcal{M}_{\psi}$, $\mathcal{V}_{\psi}$, $\mathcal{D}_{\psi}$, and $\mathcal{C}_{\psi}$ all require $O(Q_{\psi})$ operations to compute. However, these terms only need to be computed in the enriched elements, not the entire domain. Additionally, $\mathcal{M}_{\psi}$, $\mathcal{V}_{\psi}$, and $\mathcal{D}_{\psi}$ only need to be recomputed when the enrichment function changes.
@@ 6.1 Grid Resolution Study
Figures 3 and 4 show the mean velocity profiles for turbulent channel flows with $Re_{\tau}=543$, $1000$, and $2000$ with $N=10$ and $N=20$ elements respectively. The minimum distance between nodes is shown in Table 1. The results are shown for bases of order $P=5$, $7$, and $9$. The mean profiles are shifted up by $4$ for visibility as the Reynolds number is increased. For all of the cases, the mean velocity profile is in good agreement with the DNS data all the way up to the wall. There is no mismatch in the near-wall region because the no-slip boundary is enforced. The $Re_{\tau}=2000$ with $N=10$ shows some log layer mismatch, but as $N$ is increased to $20$, good agreement is seen. As expected, better agreement with the DNS data is seen for the lower Reynolds number cases and cases with higher resolution, but good agreement of the mean velocity profile is seen throughout.
@@
Figures 5 and 6 show the rms quantities for these cases with offsets for visibility. The rms data show similar trends to the DNS data. For the more underresolved cases in terms of polynomial order and Reynolds number, the rms values are underpredicted. However, as $P$ and $N$ are increased, better agreement is seen. In particular, the $N=20$, $P=9$ cases for $Re_{\tau}=543$ and $1000$ show good agreement in the rms data. More than the mean velocity profile, the rms data approaches the DNS data with increased resolution. This indicates that the wall-model is helping the mean velocity profile match the desired flow conditions, while the polynomial SEM solution is capturing the second-order turbulent statistics outside of the boundary layer.
@@
In the more underresolved cases, cusps in the rms data can be seen at the end of the first element. For example, for $Re_{\tau}=2000$, rms data cusps can be seen at $y^{+}=200$. This location corresponds with the end of the wall-adjacent element, which is the matching point location. This behavior is a result of the discontinuous gradients because only $C^{0}$ continuity is enforced at element boundaries. It is not specific to the enrichment method, but is commonly seen in under-resolved SEM simulations [34] [35] [36]. It is a sign of an under-resolved mesh [34], which is expected because the large wall-adjacent element motivates the need for the wall model. The cusps converge to zero as the resolution increases, as can be seen for an increase in the polynomial in Fig 6. Despite this inaccuracy in the gradients, good agreement in the mean velocity profile is seen, highlighting the strength of the enrichment wall-model. Additionally, Figures 5 and 6 show that this can be remedied with increased resolution, like increasing $P$.
Table 2 lists the size of the wall-adjacent element and the computed friction velocity for the channel flow tests with $Re_{\tau}=543$, $1000$, and $2000$. The wall-adjacent elements are hundreds of plus units long and extend well into the bulk of the flow. The enrichment wall-model enables the use of very large elements. The resolutions shown are far from the WRLES limit, so strict convergence of $u_{\tau}$ is not expected. However, the computed friction velocities show good agreement with the DNS data with errors under 4%. Hence, the WMLES results are modeling the correct flow behavior.
@@
For all three $Re_{\tau}$ values, increasing $P$ improves the mean velocity profile and rms quantities. This shows that the model still allows the solution to converge toward the DNS with increasing resolution as it would in the traditional SEM. Because the form of the wall-model does not change with increased $N$ or $P$, the convergence to the DNS profile is not expected to be at the order of the scheme. However, the convergence of the scheme shows that the wall-model is effective across different resolutions and that increasing resolution improves the results as expected. Table 2 shows that large elements are being used near the wall in all cases. For all three Reynolds numbers, resolutions with the first element around $y_{1}^{+}=200$ show good agreement with the mean but under-predicted rms quantities. However, for $y_{1}^{+}\leq 100$ both the mean and rms quantities show good agreement with the DNS data. While both lead to convergence, changing the size of the elements leads to more improvement, especially for the turbulent fluctuations. Increasing the Reynolds number requires additional refinement of $N$ or $P$ because of the larger gradients in the boundary layer profile.
@@
Case $\Delta_{x}^{+}$ $\Delta_{y}^{+}$ $\Delta_{z}^{+}$ $\Delta_{x}^{+}$ $\Delta_{y}^{+}$ $\Delta_{z}^{+}$ $\Delta_{x}^{+}$ $\Delta_{y}^{+}$ $\Delta_{z}^{+}$
N=10 P=5 40.4 12.9 20.2 76.4 24.0 38.2 154. 49.0 77.0
N=10 P=7 22.0 7.01 11.0 41.1 13.1 20.5 83.3 26.5 41.6
N=10 P=9 14.1 4.49 7.04 25.6 8.16 12.8 51.8 16.5 25.9
N=20 P=5 19.7 6.30 9.83 36.8 11.7 18.4 75.1 23.9 37.6
N=20 P=7 11.0 3.50 5.5 20.3 6.45 10.1 41.1 13.1 20.6
N=20 P=9 6.92 2.20 3.46 12.9 4.12 6.47 25.9 8.23 12.9
Table 1: Minimum distance between nodes in each direction, $\Delta^{+}$, for WMLES with the enrichment wall-model.
@@
Figure 5: Rms from WMLES of turbulent channel flow with $P=5,7,9$ order basis polynomials. Results for $Re_{\tau}=543$, $Re_{\tau}=1000$, and $Re_{\tau}=2000$, with $N=10$ are shown from bottom to top, shifted up by 4 for $u^{\prime+}$ and 1 for the other quantities for visibility. The solutions use a law-of-the-wall enrichment function in the wall-adjacent elements as the wall-model. The WMLES results are compared with the DNS data from [33]. The results are compared with the DNS data from [33].
Figure 6: Rms from WMLES of turbulent channel flow with $P=5,7,9$ order basis polynomials. Results for $Re_{\tau}=543$, $Re_{\tau}=1000$, and $Re_{\tau}=2000$, with $N=20$ are shown from bottom to top, shifted up by 4 for $u^{\prime+}$ and 1 for the other quantities for visibility. The solutions use a law-of-the-wall enrichment function in the wall-adjacent elements as the wall-model. The WMLES results are compared with the DNS data from [33]. The results are compared with the DNS data from [33].
$Re_{\tau}=543$ $Re_{\tau}=1000$ $Re_{\tau}=2000$
Case $y_{1}^{+}$ $u_{\tau}$ $y_{1}^{+}$ $u_{\tau}$ $y_{1}^{+}$ $u_{\tau}$
DNS [33] - 0.0543 - $0$ 0.0459
N=10 P=5 109.55 0.0548 206.89 0.0517 417.06 0.0480
N=10 P=7 109.33 0.0547 203.77 0.0509 413.46 0.0475
N=10 P=9 111.32 0.0557 202.72 0.0507 409.46 0.0471
N=20 P=5 54.28 0.0533 99.80 0.0499 203.51 0.0468
N=20 P=7 54.60 0.0546 100.63 0.0503 204.09 0.0469
N=20 P=9 54.72 0.0547 102.37 0.0512 204.60 0.0471
Table 2: Friction velocity, $u_{\tau}$, and wall-adjacent element size, $y_{1}^{+}$, for WMLES with the enrichment wall-model.
@@ 6.3 Comparison with Shear Stress Wall-Model
In this section, the enrichment wall-model is compared with purely polynomial standard SEM and a shear stress wall-modeled SEM to show the benefits. All three methods use the same initial conditions and problem specifications from Sec. 6. The standard SEM uses no-slip walls and no added filtering, so the numerics match the enriched case exactly. At the coarse resolutions in the near-wall region that will be presented, using a filter for added dissipation in the SEM leads to large oscillations and a mismatch of the targeted turbulent profile. The shear stress wall-model is implemented as described earlier in [4] [37]. The shear stress wall-model uses no-penetration slip-wall boundary conditions that apply a modeled shear stress at the wall. The matching point is taken as the top of the first wall-adjacent element. The model uses the Reichardt law-of-the-wall as the equilibrium wall-stress model. The initial solution is allowed to develop for $8$ flow-through times before the shear stress wall-model is activated. Additionally, an explicit filter with weight $0.05$ and cutoff ratio $0.65$ are used to add dissipation and greatly decrease the amount of slip at the wall. This added dissipation serves as the subgrid scale model for this LES, as is often done in SEM [35] [38].

### v3

@@ 1 Introduction
This study seeks to make use of the high-order polynomial solution of SEM to develop a different kind of wall-model better suited for high-order methods than traditional models. The wall-model builds upon prior work on the development of so-called enriched basis turbulence wall models in the context of DG [8] [13] and the continuous finite element (FEM) [14] [15] [16]. These methods leverage the polynomial solution and augment it by considering the empirical wall function as part of the solution representation. Doing so prevents spurious oscillations in the near-wall region due to insufficient polynomial resolution of the boundary layer and improves the solution representation by allowing the degrees of freedom to represent the flow fluctuations instead of the mean flow behavior. This work serves to develop the framework for enrichment wall-models in SEM so that SEM because of its low dissipation and dispersion errors for its cost [9] and its computational efficiency and scalability [10].
@@ 2 Governing Equations
where $\bm{u}$ is the velocity vector, $t$ corresponds to time, $p$ is the pressure, and $\nu$ is kinematic viscosity of the fluid. The equations are solved over a computational domain $\Omega$ with boundary $\partial\Omega$. In this work, we consider implicit LES (iLES), where the numerical dissipation of the scheme is used to represent the small scales, instead of an explicit subgrid-scale turbulence model.
@@ 3 Spectral Element Method
Assuming $\bm{u}^{n-1}$ is $C^{0}$ continuous, $\bm{u}^{n}$, will be $C^{0}$ continuous if $\Delta\bm{u}^{n}$ is also $C^{0}$ continuous. Hence, because the operators in Eq. (21)-(22) enforce $C^{0}$ continuity of $\Delta\bm{u}^{n}$, $\bm{u}^{n}$ will maintain the appropriate continuity at all timesteps.
@@ 4 Enrichment Method
The enriched SEM begins by putting the governing equations into weighted residual form as shown in Eqs. 3-4. However, for the enriched SEM, the solution is $(\bm{u},p)\in X_{b,\psi}^{P}(\Omega)\times Y^{P}(\Omega)$, where $X_{\psi}\subset H^{1}(\Omega)$ such that $X_{\psi}^{P}=\{u\,|\,u=u_{p}+\psi,u_{p}\in X^{P}(\Omega_{\psi})\cap X^{P}(\Omega_{p}),\psi\in H^{1}(\Omega_{\psi})\cap H^{1}(\Omega_{p}),\psi(\bm{x})=0\,\forall\bm{x}\in\Omega_{p}\}$, where $\Omega_{\psi}$ and $\Omega_{p}$ are the enriched and unenriched parts of the domain respectively, such that $\Omega_{\psi}\cup\Omega_{p}=\Omega$ and $\Omega_{\psi}$ and $\Omega_{p}$ do not overlap. $X_{b,\psi}$ is the subset of $X_{\psi}$ such that Dirichlet boundary conditions are satisfied on the boundary $\partial\Omega$. As is done in the traditional SEM, the weighted residual form is integrated by parts to give Eqs. 5-6. From here, the enriched solution representation from Eq. 24 is substituted into the weighted residual form:
@@
For the application to WMLES, we assume the enrichment function is constant within a timestep and is adapted explicitly using matching points, which will be detailed in Sec. 5. Hence, the time derivative of the enrichment function is put to the right hand side of Eq. (28). With the governing equations in this form, the left-hand side and the first three terms on the right-hand side are the same as in the traditional SEM and the last five terms on the right-hand side are additions from the enrichment [31]. This separability allows the presented enrichment method can be easily implemented with the addition of associated terms into existing solvers as right-hand side terms.
@@
Because the enrichment function is constant within each timestep, the change in the enriched solution is equal to the change in the polynomial component: $\Delta\bm{u}^{n}=\Delta\bm{u}^{n}_{p}$. Further, Lemma 4.1 proves that $\Delta u_{p}^{n}\in H^{1}(\Omega)$. Hence, the standard SEM scatter-gather operations ensure that the solution is continuous. [27]:
@@
Lemma 4.1.
If $\psi^{n}$ is continuous within a timestep, the change in the solution, $\Delta u_{p}^{n}$ is in $X^{P}(\Omega)$.
Proof.
Let $u_{p,\psi}$ and $u_{p,p}$ be the polynomial portions of the solution in $\Omega_{\psi}$ and $\Omega_{p}$ respectively. $u_{p,\psi}^{n}$ and $u_{p,\psi}^{n-1}$ are in $X^{P}(\Omega_{\psi})$. Hence $\Delta u_{p,\psi}^{n}\in X^{P}(\Omega_{\psi})$. Similarly, $\Delta u_{p,p}^{n}\in X^{P}(\Omega_{p})$. Let $\Gamma_{\psi,p}$ be the boundary between $\Omega_{\psi}$ and $\Omega_{p}$. Within $\Omega_{\psi}$, the value of $u(\Gamma_{\psi,p})=u_{p,\psi}(\Gamma_{\psi,p})+\psi(\Gamma_{\psi,p})$. Hence, because $\psi^{n}$ is constant within the timestep,
$\Delta u^{n}(\Gamma_{\psi,p})$ $=u^{n}_{p,\psi}(\Gamma_{\psi,p})+\psi^{n}(\Gamma_{\psi,p})-u^{n-1}_{p,\psi}(\Gamma_{\psi,p})-\psi^{n}(\Gamma_{\psi,p}),$ (44)
$=u^{n}_{p,\psi}(\Gamma_{\psi,p})-u^{n-1}_{p,\psi}(\Gamma_{\psi,p}),$ (45)
$=\Delta u^{n}_{p,\psi}.$ (46)
Similarly, within $\Omega_{p}$, $u(\Gamma_{\psi,p})=u_{p,p}(\Gamma_{\psi,p})$, so
$\Delta u^{n}(\Gamma_{\psi,p})$ $=u^{n}_{p,p}(\Gamma_{\psi,p})-u^{n-1}_{p,p}(\Gamma_{\psi,p}),$ (47)
$=\Delta u^{n}_{p,p}.$ (48)
Because $u^{n},u^{n-1}\in H^{1}(\Omega)$, $\Delta u^{n}(\Gamma_{\psi,p})$ is equal between both parts of the domain. Hence $\Delta u^{n}_{p,\psi}(\Gamma_{\psi,p})=\Delta u^{n}_{p,p}(\Gamma_{\psi,p})$. $\Delta u_{p}^{n}$ is in $X^{P}(\Omega_{\psi})\cap X^{P}(\Omega_{p})$ and is continuous at $\Omega_{\psi}\cap\Omega_{p}$. Hence $\Delta u_{p}^{n}\in X^{P}(\Omega)$. ∎
@@ 6.1 Grid Resolution Study
Figures 3 - 6 show the mean velocity and rms profiles for turbulent channel flows with $Re_{\tau}=543$, $1000$, and $2000$ with $N=10$ and $N=20$ elements respectively. Table 1 lists the size of the wall-adjacent element for the channel flow tests with $Re_{\tau}=543$, $1000$, and $2000$. The wall-adjacent elements are hundreds of plus units long and extend well into the bulk of the flow. The use of such large elements is enabled by the the enrichment wall-model. The results are shown for bases of order $P=5$, $7$, and $9$. The mean profiles are shifted up by $4$ for visibility as the Reynolds number is increased. For all of the cases, the mean velocity profile is in good agreement with the DNS data all the way up to the wall. There is no mismatch in the near-wall region because the no-slip boundary is enforced. The $Re_{\tau}=2000$ with $N=10$ shows some log layer mismatch, but as $N$ is increased to $20$, good agreement is seen. The rms data show similar trends to the DNS data. For the more underresolved cases, the rms values are underpredicted, however, as $P$ and $N$ are increased, better agreement is seen. In particular, the $N=20$, $P=9$ cases for $Re_{\tau}=543$ and $1000$ show good agreement in the rms data. As expected, better agreement with the DNS data is seen for the lower Reynolds number cases and cases with higher resolution. This indicates that the wall-model is helping the mean velocity profile match the desired flow conditions, while the polynomial SEM solution is capturing the second-order turbulent statistics outside of the boundary layer.
@@
In the more underresolved cases, cusps in the rms data can be seen at the end of the first element. For example, for $Re_{\tau}=2000$, rms data cusps can be seen at $y^{+}=200$. This location corresponds with the end of the wall-adjacent element, which is the matching point location. This behavior is a result of the discontinuous gradients because only $C^{0}$ continuity is enforced at element boundaries. It is not specific to the enrichment method, but is commonly seen in under-resolved SEM simulations [34] [35] [36]. It is a sign of an under-resolved mesh [34], which is expected because the large wall-adjacent element motivates the need for the wall model. Figures 5 and 6 show that this can be remedied with increased resolution, like increasing $P$. Despite this inaccuracy in the gradients, good agreement in the mean velocity profile is seen, highlighting the strength of the enrichment wall-model.
@@
For all three $Re_{\tau}$ values, increasing $P$ and $N$ improves the agreement with the DNS as expected. Table 1 shows that large elements are being used near the wall in all cases. For all three Reynolds numbers, resolutions with the first element around $y_{1}^{+}=200$ show good agreement with the mean but under-predicted rms quantities. However, for $y_{1}^{+}\leq 100$ both the mean and rms quantities show good agreement with the DNS data. Hence, refining $N$ leads to more improvement than refining $P$, especially for the turbulent fluctuations. The larger Reynolds number cases require more refinement than the lower because of the larger gradients in the boundary layer profile. Overall, the enrichment wall-model effectively represents the near-wall behavior of the turbulent flow with large elements for a range of Reynolds numbers.
@@
Case $y_{1}^{+}$ $y_{1}^{+}$ $y_{1}^{+}$
N=10 P=5 109.55 206.89 417.06
N=10 P=7 109.33 203.77 413.46
N=10 P=9 111.32 202.72 409.46
N=20 P=5 54.28 99.80 203.51
N=20 P=7 54.60 100.63 204.09
N=20 P=9 54.72 102.37 204.60
Table 1: Wall-adjacent element size, $y_{1}^{+}$, for WMLES with the enrichment wall-model.
@@
Figure 5: Rms from WMLES of turbulent channel flow with $P=5,7,9$ order basis polynomials. Results for $Re_{\tau}=543$, $Re_{\tau}=1000$, and $Re_{\tau}=2000$, with $N=10$ are shown from bottom to top, shifted up by 4 for $u^{\prime+}$ and 1 for the other quantities for visibility. The solutions use a law-of-the-wall enrichment function in the wall-adjacent elements as the wall-model. The WMLES results are compared with the DNS data from [33].
Figure 6: Rms from WMLES of turbulent channel flow with $P=5,7,9$ order basis polynomials. Results for $Re_{\tau}=543$, $Re_{\tau}=1000$, and $Re_{\tau}=2000$, with $N=20$ are shown from bottom to top, shifted up by 4 for $u^{\prime+}$ and 1 for the other quantities for visibility. The solutions use a law-of-the-wall enrichment function in the wall-adjacent elements as the wall-model. The WMLES results are compared with the DNS data from [33].
@@ 6.3 Comparison with Shear Stress Wall-Model
In this section, the enrichment wall-model is compared with purely polynomial standard SEM and a shear stress wall-modeled SEM to show the benefits. All three methods use the same initial conditions and problem specifications from Sec. 6. The standard SEM uses no-slip walls and no added filtering, so the numerics match the enriched case exactly. At the coarse near-wall resolutions presented, using a filter for added dissipation in the SEM leads to large oscillations and a mismatch of the targeted turbulent profile. The shear stress wall-model is implemented as described earlier in [4] [37]. The shear stress wall-model uses no-penetration slip-wall boundary conditions that apply a modeled shear stress at the wall. The matching point is taken as the top of the first wall-adjacent element. The model uses the Reichardt law-of-the-wall as the equilibrium wall-stress model. The initial solution is allowed to develop for $8$ flow-through times before the shear stress wall-model is activated. Additionally, an explicit filter with weight $0.05$ and cutoff ratio $0.65$ are used to add dissipation and greatly decrease the amount of slip at the wall. This added dissipation serves as the subgrid scale model for this LES, as is often done in SEM [35] [38].